Baba Software's current balance sheet shows total common equity of $50,125,000. The company has 5,010,000 shares of stock outstanding, and they sell at a price of $28.50 per share. By how much do the firm's market and book values per share differ?
Answer #1

Answer:

Given

Book value of shares=$50,125,000

Shares outstanding=5,010,000

Market value of per share=$28.50

Book value of per share=50125000/5010000=$10.005

Difference between book value and market value per share=Market value of per share-Book value of per share=28.5-10.005=$18.495